### v1

# SIGGesture: Generalized Co-Speech Gesture Synthesis via Semantic Injection with Large-Scale Pre-Training Diffusion Models

Abstract
The automated synthesis of high-quality 3D gestures from speech is of significant value in virtual humans and gaming. Previous methods focus on synthesizing gestures that are synchronized with speech rhythm, yet they frequently overlook the inclusion of semantic gestures. These are sparse and follow a long-tailed distribution across the gesture sequence, making them difficult to learn in an end-to-end manner. Moreover, generating gestures, rhythmically aligned with speech, faces a significant issue that cannot be generalized to in-the-wild speeches. To address these issues, we introduce SIGGesture, a novel diffusion-based approach for synthesizing realistic gestures that are of both high quality and semantically pertinent. Specifically, we firstly build a strong diffusion-based foundation model for rhythmical gesture synthesis by pre-training it on a collected large-scale dataset with pseudo labels. Secondly, we leverage the powerful generalization capabilities of Large Language Models (LLMs) to generate proper semantic gestures for the various speech content. Finally, we propose a semantic injection module to infuse semantic information into the synthesized results during diffusion reverse process. Extensive experiments demonstrate that the proposed SIGGesture significantly outperforms existing baselines and shows excellent generalization and controllability.
Keywords:
Co-speech gesture synthesis Semantic gestures Large Language Models Diffusion models
1 Introduction
Body gesture can greatly improve people’s expressiveness in communication, which can emphasize or convey specific semantics with non-verb body movements [8, 18, 49, 36]. Besides, high-quality 3D gesture animation can present vivid virtual characters that help listeners to concentrate and improve the intimacy between humans and virtual characters. Therefore, there is a growing need for automatic high-quality co-speech gesture synthesis in many applications, including Virtual Humans, non-player gaming characters, and robot assistants. Many researchers have proposed lots of approaches including rule-based methods [24, 35] and data-driven methods [52, 53, 2].
Figure 1: The framework of the proposed method. The upper shows the training details of diffusion process and denoising process. The lower shows the inference process of synthesizing the gesture from the conditions. Specifically, given audio feature, speaker identity, and speech content, the audio feature and speaker identity are directly fed into the denoising network. The speech content are as input in the interaction with LLM, and the LLM will generate a set of candidate semantic gestures from the input text for the next semantic injection process.
3D co-speech gesture generation aims to generate gestures with natural movements, accurate rhythm, and clear semantics based on a given speech. Rule-based methods [24, 35] achieve this goal by hard-rules, such as keyword "good" with a simple gesture "thump up". However, these approaches can only deal with simple semantic gestures, which are lack of naturalness, smoothness, and diversity. The principal reason for this is that generating co-speech gestures poses a challenge due to its inherently complex many-to-many mapping. Recently, deep learning with large-scale dataset has achieved many breakthroughs, especially in content synthesis. With the aid of deep learning, co-speech gesture synthesis also bloomed many remarkable methods. These methods attempt to solve this problem by training a conditional model with multi-modal inputs (audio, text, speaker identities).
Co-speech gestures include rhythmic gesture and semantic gesture. Some works attempt to generate semantic gestures in co-speech gesture generation, such as GestureDiffuCLIP [2], LivelySpeaker [52], and Bodyformer [38]. Both GestureDiffuCLIP [2] and LivelySpeaker [52] use CLIP [40] to learn the relationship between semantic text and motion. However, semantic gesture constitutes only a minor portion of the entire sequence and follows a long-tailed distribution, complicating the direct learning of its multi-modal relationships. Moreover, synthesizing gestures that are rhythmically synchronized with the given speech faces a challenge of generalizing to in-the-wild speech scenarios.
Table 1: Comparison of Co-Speech Gesture Datasets. We ignore some datasets with 2D and 3D key-points label. "3D Rot" indicates 3D skeleton with bone rotation. "PGT" means Pseudo Ground Truth.
Trinity
Mocap
2017 [14]
TWH
Mocap
2019 [28]
Hahibie et al.
S2G-3D
2021 [20]
BEAT(X)
Mocap
2022 [32]
ZEEG
Mocap
2022 [16]
Yoon. et al.
TED-SMPL
2023 [34]
Talkshow
S2G-SMPL
2023 [48]
LSMoG
Ours
2024
Data type 3D Rot 3D Rot 3D Rot 3D 3D Rot PGT 3D Rot PGT 3D Rot PGT 3D Rot
Duration (hours) 4 20 38 76 4 30 27 393
For addressing the after-mentioned issues, we propose a diffusion-based semantic injection method with large language model for semantic gesture synthesis, which has excellent controllability, interpretability, and generalization ability. It begins with a insight that real-word human conversation contains limited semantic gestures. Besides, for a specific semantic word, such as "wonderful", different people usually present similar gesture such as "opening both arms". During the diffusion reverse process, we can utilize semantic matched gestures as prompts to guide the generative model in synthesizing the appropriate semantic gestures. Therefore, we collect a large-scale semantic motion database, which can include most semantic motion in communication. Large Language Models (LLMs), such as GPT3 [6] and GPT-4, have shown their ability not only as proficient models for natural language processing (NLP) tasks but also as potent instruments for tackling complex problems. Therefore, we introduce large-language model to assign contextually appropriate gestures to the corresponding text. Leveraging the strong semantic analysis capabilities of LLMs, the proposed method is adept at handling various languages.
For a specific speech, different people will present notably different rhythmic gesture. For such many-to-many issue, the best solution is to use deep learning with large-scale dataset to learn their complex relationship. Therefore, it is necessary to build a robust foundational model for rhythmic gesture. Then, the semantic gesture can be injected into the rhythmic gesture with naturalness. However, the co-speech gesture generation research community lack of 3D skeletal data because of expensive motion capture system, which is different from other synthesis tasks, such as text2image, text2speech. Therefore, for improving the basic quality of rhythmic gesture, we collect a large-scale gesture dataset named LSMoG for pre-training that reaches to 400 hours. Table 1 shows the statistical data of co-speech gesture datasets. Some datasets with 2D and 3D key-points label are ignored because they are difficult to use in real applications. To the best of our knowledge, it is the largest dataset in the research of co-speech gesture synthesis field. By the optimization of pre-traing and fine-tuning, the synthesized gesture motion has excellent naturalness and generalization.
The main contributions of our works contains the following aspects:
  1. 1.
    We propose a LLMs-based semantic augmentation method for co-speech gesture synthesis, which can greatly improve the generalization ability in semantic gesture synthesis. For controllable semantic gesture synthesis, we build a set of semantic gesture clips and use LLMs to generate proper candidate gestures for a specific speech content.
  2. 2.
    We build a strong diffusion-based foundation model for rhythmical gesture synthesis by pre-training it on a collected large-scale dataset with pseudo labels. This dataset is the largest dataset for co-speech gesture synthesis, which contains approximately 400 hours of motion sequences.
  3. 3.
    Extensive experiments show that the proposed method outperforms state-of-the-art methods by a large margin. Especially, the visualization comparison indicates that our method has more stable, expressive, and robust results than other approaches.
2 Related Work
2.1 Co-Speech Gesture Generation
The research of co-speech gesture generation can be divided into two branches, Rule-based methods [9, 35, 10] and learning-based methods [1, 2, 33, 17, 33]. Rule-based works [9, 35, 10] for co-speech gesture synthesis are mainly generated from a pre-defined motion database by keyword matching or other specific rules. These rule-based based methods require lots of human effort in defining gesture units and complex mapping rules, which is costly and inefficient. Besides, the results of rule-based methods are usually lack of smoothness and naturalness.
Co-speech gesture synthesis is a complex problem that requires consideration of the audio, the gesture, and their many-to-many relationships. Benefit from recent advanced developments of deep learning, all of recent methods [1, 2, 33] adopt deep neural networks as a strong tool to learn the complex relationships from audio to gesture in an end-to-end manner. Earlier works [17, 39] explore various network architecture to regress 2D keypoints of the human body movements because of lacking high-quality 3D gesture datasets. These 2D datasets are constructed by using off-the-shelf pose estimator [7] to obtain pseudo 2D gesture annotations from the online videos. However, these 2d results are lack of expressiveness in real application. Recently, some high-quality speech-gesture corpus such as BEAT [32] and ZeroEGGS [16] are released, which contribute to the development of co-speech gesture synthesis. Existing works focus on exploring the effectiveness of different network architectures, including the multi-layer perceptron (MLP) [27], convolutional neural networks (CNN) [20, 48], recurrent neural networks (RNNs) [49, 22, 33, 3] and Transformers [4, 38]. Besides the exploration of model architecture, some approaches [17, 44] dive into researching the connections between co-speech gesture and speech audio, text transcript, speaking style, and speaker identity. Besides, some approaches involve the adversarial training [17, 33], Phase-Guided Motion Matching[45], and reinforcement learning [42] to guarantee realistic results. In addition to gestures, some methods [48, 11, 31] begin to explore the generation of full-body movements, including body gestures and facial expressions.
The rapid development of diffusion models [23, 41] has surpassed the traditional synthesis paradigm such as GAN [19] and VAE [26], showcasing a powerful and realistic characteristic in synthesis. Diffusion-based motion generation also become a popular research direction, such as text2motion [50]. To be specific, text-guided motion can synthesize realistic and expressive results with diffusion model. In co-speech gesture synthesis, recent approaches such as GestureDiffuCLIP[2], LivelySpeaker [52], DiffStyleGesture [44], C2G2[25] are diffusion-based approaches.
2.2 Semantic-aware Co-Speech Gesture Generation
In gesture language, semantic gesture units are crucial for conveying specific intentions, thoughts, emotions, and enhancing gestural expressiveness. For generating more expressive gestures, there are also some attempts such as Gesticulator [27], GestureDiffuCLIP[2], LivelySpeaker[52], Bodyformer [38]. Specifically, GestureDiffuCLIP[2] uses CLIP [21] to align motion and text semantic, then use the text as additional condition to control the semantic gestures. LivelySpeaker[52] is a two-stage framework for semantic and rhythmic gesture generation, which aim at decoupling the semantic gesture generation and rhythmic gesture generation. In semantic gesture generation, they also use a CLIP style module to align the semantic motion and semantic text. Some approaches add semantic text as additional input to fuse semantic into the model, such as Rhythmic Gesticulator [1], Gesticulator [27], FreeTalker [46] and Trimodal [49]. BodyFormer [38] takes both low-level and high-level speech features as input and generates a sequence of realistic 3D body gestures in an auto-regressive manner. The high-level speech features represent the semantic gestures. With the development of LLMs, some approaches also use it to extract motion-related information from textual input, such as GesGPT[15], MoConVQ [47].
3 Method
3.1 Problem Formulation
Given a specific audio, its corresponding text, and identity information, the proposed method can generate high-quality 3D skeletal co-speech gesture results. To be specific, considering a sample of raw audio ${A=\{a_{0},a_{1},...,a_{N}\}}$, we adopt Jukebox [13] to extract its acoustic feature, which is trained on large-scale music datasets. All of the motion sequences =$\mathcal{M}=\{m_{0},m_{1},...,m_{T}\}$ are encoded into discrete tokens by VQVAE[37]. With VQVAE, a motion can be represented by latent embedding $\hat{e}=\{e_{0},e_{1},...,e_{T}\}$. Then, we adopt transformer-based diffusion model as the denoising network to recover the latent embedding by removing the noise. The reverse denoising process of the optimized diffusion model $G$ parameterized by $\theta$ can recover the motion latent embedding $\hat{e}=\{e_{0},e_{1},...,e_{T}\}$ conditioned on the speech audio sequence and identity information. With integration with LLM, it can provide accurate candidate semantic gestures for the speech content. Finally, the candidate semantic gesture embedding are fused into the synthesized motion latent embedding by semantic injection module.
3.2 Semantic Gesture Collection
Through extensive observation of numerous instances, it has been discerned that semantic gestures maintain a universality in gestural communication. Specifically, it is evident that different individuals display similar gestural expressions when articulating identical semantic concepts. This observation provides a feasible methodology for the collection of an extensive array of semantic gesture actions, aimed at covering a comprehensive range of semantic contexts. We collect a large-scale semantic motion clip $G=\{G_{0},G_{1},...,G_{N}\}$ and text $T=\{T_{0},T_{1},...,T_{N}\}$ pairs from BEAT dataset by manual selection. $N$ is the number of the semantic gestures. All of the semantic clip is shorter than 3 seconds and each motion clip represent a whole semantic motion. After selection, we have collected 2,537 semantic motion clips. Some examples of semantic gestures are shown in Figure 2. These motion clips are sorted into a predefined motion semantic system.
Figure 2: The statistical data of semantic gesture dataset (upper left part), some examples of semantic label (lower left part), and some semantic gesture examples (right part).
All of semantic gesture clips are encoded into latent embedding $V=\{v_{0},v_{1},...,v_{N}\}$ by a pre-trained VQVAE model. To improve the diversity of semantic gesture, we add a random noise to the embeddings during semantic injection.
3.3 Diffusion models with semantic injection
SIGGesture employs a diffusion model to synthesize the latent representation of gesture, which consists of a diffusion process and a denoising process, as shown in upper part of Figure 1. Given a specific real motion data distribution $p(x_{0})$, our goal is learn a distribution $q_{\theta}(x_{0})$ that is parameterized by ${\theta}$.
Diffusion process. We follow the previous DDPM [23] definition of diffusion as a Markov noising process with latents $x_{0:T}$ that follow a forward noising process $p(x_{0:T}|x_{0})$. Here, $x_{0}$ is the latent motion data and same as previous notation $\hat{e}$. The forward diffusion process is defined as a Markov chain that gradually adds Gaussian noise to the latent representation of motion, as follows
$\begin{split}q(x_{0:T}|x_{0})=\prod_{t=1}^{T}q(x_{t}|x_{t-1}),\end{split}$ (1)
where
$\begin{split}q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}{I}).\end{split}$ (2)
$\beta_{t}$ are variance schedule hyperparameters that controls the distribution of Gaussian noise. In the reverse denoising process, $\beta_{t}$ are pre-defined constant parameters. By this diffusion process, the original latent representation of motion $x_{0}$ progressively substituted by random noises. When $T$ approaches infinity, the distribution of $x_{T}$ follows a pure white noise distribution.
Denoising process. Through diffusion process, the real data are transferred into pure white noise by adding random noise. On the contrary, the denoising process recovers the real data from pure white noise by removing the random noise. The unconditional reverse process of recover a gesture motion can be represented by the following formula
$\begin{split}p_{\theta}(x_{0:T})=p_{\theta}(x_{0})\prod_{t=1}^{T}p_{\theta}(x_{t-1}|x_{t}),\end{split}$ (3)
where
$\begin{split}p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t};u_{\theta}(x_{t},t),\beta_{t}I).\end{split}$ (4)
The noise $x_{t}$ is sampled from the following defined distribution
$\begin{split}p(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha_{t}}},(1-\bar{\alpha_{t}}){I}).\end{split}$ (5)
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha_{t}}=\prod_{k=0}^{t}\alpha_{k}$. The above explanation shows the process of unconditioned diffusion generation. For conditional content synthesis, the conditions should be injected into the diffusion generation procedure. The details of the denoising network are shown in the upper right part of Figure 1. In our method, the conditions are audio $c_{audio}$ and speaker identity $c_{id}$ respectively. For simple notation, we mark the two conditions as $c=[c_{audio},c_{id}]$. It should be noted that the audio feature is extracted by Jukebox [13] instead of WavLM [12]. Jukebox [13], trained on large-scale music datasets, can focus on the rhythmic features rather than the speech semantic content. By injecting the condition $c$ into the generation process, the reverse process of each time step can be updated by the following formula
$\begin{split}p_{\theta}(x_{t-1}|x_{t},c)=\mathcal{N}(x_{t};u_{\theta}(x_{t},t,c),\beta_{t}I).\end{split}$ (6)
In a short summary, we firstly start denosing process by sampling $x_{T}$ from a pure white nose distribution $\mathcal{N}(0,I)$. Then, we iteratively denoise the latent variable $x_{t}$ to obtain the final results $x_{0}$.
Training objective. For a speech condition $c$, we reverse the forward diffusion process by learning to estimate $p(x_{t},t,c)$ approximate $x_{0}$ with a parameterized model for all times step $t$. We use a loss function proposed in [23] to optimize the whole model, as following,
$\begin{split}\mathcal{L}={E_{x,t}[\parallel x-p(x_{t},t,c)\parallel^{2}_{2}]}.\end{split}$ (7)
During training, the model is trained under conditional and unconditional manner with a specific probability. Once the training converges, the the diffusion model can predict the noise parameters by considering the unconditional setting and condition $c$.
Inference with semantic injection. The inference process involves solely the denoising process of diffusion models. The pipeline of inference process is shown in the lower part of Figure 1. To be specific, speech with its corresponding text will be feed into the generation pipeline. It should be noted that the words in the text is aligned with speech in timeline. Each words have a duration time. Then, the text are fed into the LLMs to produce candidate semantic gestures $\mathcal{G}=(G_{0},...,G_{t})$ for the speech content. The interaction with LLMs process is shown in the left-top corner of inference pipeline in Figure 1. With the ability of LLM, the proposed method can process various language, such as Chinese and Japanese.
The candidate semantic gesture is encoded into latent embedding $\mathcal{P}=(p_{0},...,p_{n})$, where $n$ is number of candidate semantic gestures. We set a control $K$ to determine the degree of semantic injection. From time step $T$ to $K$, the denoising latent variable is marked as $\hat{x_{t}}$ that contains the semantic information.
$\begin{split}p_{\theta}(x_{K:T})=p_{\theta}(x_{T})\prod_{t=K}^{T}p_{\theta}(x_{t-1}|\hat{x}_{t}).\end{split}$ (8)
The semantic gesture $P$ can be injected into the latent variable $x_{t}$ by the follow formula
$\begin{split}\hat{x}_{t}=x_{t}*m+P*(1-m),\end{split}$ (9)
where $m$ denotes the timeline mask of semantic words. From time step $K$ to 1, the denoising latent variable $x_{t}$ is not with semantic injection, which can preserve the structure and smoothness of the generated gesture.
Generating long sequence. For synthesizing long motion sequence, the common practice is autoregressive method, which generates content sequentially. However, generative models trained on small pieces may cause unnaturalness and defectiveness. Instead of auto-regressive manner, we adopt DiffCollage [51] to generate long motion sequence.
3.4 Large-scale motion corpus collection
In 3D animation, high-quality data collection is remarkably expensive, professional, and time-consuming. High-quality motion capture requires specialized cameras, suits, and markers. Besides, it often requires experienced technicians, actors, and animators to set up, perform, and process the captured data. The process of capturing, cleaning, and integrating motion data is labor-intensive and time-consuming, contributing to the overall expense. In recent years, several high-quality datasets are released to the research community, as shown in Table 1. BEAT [32] is the largest dataset for co-speech gesture synthesis, which contains 76 hours data. However, data-driven approaches with larger models are also short for data. In the inspiration of large-language training paradigm, noise data for pre-training and high-quality data for fine-tuning, we aim at collecting a large-scale gesture corpus for pre-training by video motion capture technique such as HybridIK [29], SMPLify [5]. Before running the video motion capture approach, we collect a large mount of videos about speech or talk from the Internet, such as TED. Then, we conduct the data cleaning pipeline to obtain the final shots of interests. Table 2 shows the statistical data of the motion corpus collection. Finally, we obtain about 400 hours 3D gesture assets of speech. As shown in Table 1, the proposed dataset is significantly larger than previous datasets by at least 6 times.
Table 2: Statistical data of the motion corpus collection.
Items Number Average length Total length
original videos 6,032 $\sim$23min $\sim$2,312h
shots-of-interests 138,209 $\sim$20s $\sim$746h
final results 70,784 $\sim$20s $\sim$393h
4 Experiments
4.1 Datasets
In this paper, we use the largest high-quality speech-gesture dataset BEAT to evaluate the proposed method. BEAT is constructed by using commercial motion capture system, including face emotion and body motion. It contains about 76 hours motion-audio paired data of 30 speakers talking about different topics. We deeply clean the dataset by removing some defective motion sequences such as speaker drinking, wandering, and incorrect motion. After cleaning the raw data, the BEAT dataset contains 46 hours motion data and its corresponding text and audio. During training the VQVAE model, the original Euler angles are converted to rotation matrix for better convergence. During training the diffusion-based generation model, all motions are represented by latent embeddings in VQVAE latent space. The datasets collected from Internet by video motion capture are only used for pre-training because of its relatively poor quality.
4.2 Evaluation Metrics
For content generation tasks, human objective evaluation is the most important evaluation method because of lacking clear ground and explicit criteria.Therefore, the human objective evaluation is the main evaluation method in our method. Each motion slice is re-targeted to Mixiamo model and rendered as a motion video sequence. They are asked to rate the slices from the following four aspects respectively: (1) Naturalness, (2) Rhythm, (3) Diversity, (4) Semantic. The scores assigned to each rating are in the range of 1-10, corresponding from worst to best. All scores are normalized to 0-1. Besides, we also calculate some quantitative metrics adopted by previous works. FGD [49] measures the distribution difference between generated data and ground truth. Beat Consistency (BC) [30] calculates the average distance between every audio beat and its nearest motion beat. Diversity is a metric of measuring the variations of the synthesized gesture. Semantic-Relevant Gesture Recall (SRGR) [32] measures the semantic correct key point of synthesized data by comparing it with the ground truth data.
4.3 Implementation Details
The whole training process consists of two steps. In the first step, we train the VQVAE model using the Adam optimizer for 400 epochs. In the second step, we train the diffusion-based gesture generation model using the Adam optimizer for 3000 epochs with batchsize 256. The diffusion steps for training and inference are 1000. During the training of diffusion model, we firstly use the collected data for pre-training (2000 epochs with batchsize 256), and then use the high-quality dataset for fine-tuning (1000 epochs). The length of the input audio feature is 12 seconds. We train the diffusion model on 32 V100 GPUs.
4.4 Comparing Baselines
We compare our method on BEAT dataset with the state-of-the-art methods in the last year, including DiffuseStyleGesture[44], TalkSHOW[48], QPGesture [45], EMAGE [31]. DiffuseStyleGesture[44] is a diffusion model-based co-speech gesture synthesis approach, which takes the motion style into generation process. TalkSHOW[48] uses a simple full-body speech-to-motion generation framework with auto-regressive manner, in which the face, body, and hands are modeled separately. QPGesture [45] is a quantization-based and phase-guided motion-matching approach for co-speech gesture synthesis. EMAGE [31] leverages a masked gesture reconstruction to enhance audio-conditioned gesture generation.
4.5 Evaluation Results
Figure 3: Visualization results between the proposed method and other state-of-the-art methods.
Qualitative Results For this task, there is lack of absolutely objective metric to evaluate the performance of various methods. Therefore, We strongly recommend the readers to view our demonstration video to gain an intuitive understanding of the qualitative outcomes. In Figure 3, it shows some synthesized results of the baselines conditioned on the same speech. From these results, we can observe that the synthesized co-speech gestures of the proposed method are more realistic, agile, and diverse than those of the baselines on BEAT datasets. The baselines suffer different extents of jittering in the results, especially lacking of semantic gesture and naturalness. DiffuseStyleGesture [44] is also a diffusion-based method, which directly learn the origin motion representations. However,its results lack smoothness and naturalness (e.g. foot sliding). QPGesture [45] exhibits noticeably slower and less varied motion, while ours demonstrates agility comparable to actual movement. For the specific semantic expressions, apart from our approach, other methods cannot effectively synthesize proper semantic gestures. More results are shown in Figure 4. The results indicate that our method can generate proper semantic gestures for the speech semantics. For example, "free will" refers to the first row with open arms; "left" and "right" indicate pointing to the corresponding direction. These results show the surpassing capability of SIGGesture in synthesizing semantic accurate and natural co-speech gesture.
Figure 4: The visualization results of the proposed method. The gestures in the red box are semantic gesture, which is labeled by the content in green box.
User study. As we all known, the evaluation of generative task is very subjective. Although some metrics are introduced to evaluate the performance, there is a huge gap between metric evaluation and human visual perception. We conduct a user study to evaluate the performance of different baselines. Specifically, we randomly select 50 synthesized samples and shuffle results of different methods. Following previous works[52, 44], the participants are asked to scoring the gesture based on naturalness, rhythm appropriateness, diversity, and semantic consistence for the shuffled visual results.. It’s important to note that, for diversity, we enable users to score motion diversity under the condition of smooth and natural motion. The results of user study are shown in Table 3. The results show that the generated gestures of SIGGesture are dominantly preferred on four metrics over the comparing baselines. Especially, SIGGesture exceeds these baslines by a large margin. It can be concluded that large-scale pre-training benefits rhythm and diversity, while semantic injection is advantageous for semantic consistency. Overall, SIGGesture is capable of generating more realistic, synchronized, diverse, and understandable gestures that humans prefer.
Table 3: The statistical data of user study by comparing the proposed method to baselines. The $\downarrow$" means the lower, the better. The $\uparrow$" means the higher, the better.
[…]
DiffuseStyleGesture (IJCAI’23) [44] 0.78 0.82 0.65 0.57
TalkSHOW (CVPR’23)[48] 0.56 0.63 0.57 0.37
QPGesture (CVPR’23)[45] 0.62 0.72 0.53 0.56
EMAGE (24) [31] 0.80 0.81 0.79 0.60
SIGGesture 0.82 0.87 0.83 0.92
Quantitative Results. We compare our method with the baselines using four metrics on BEAT dataset. The comparison results are shown in Table 4. As discussed in recent works [43, 2], these metrics have obvious weakness because these value can not keep consistent with their visual results. For example, TalkSHOW[48] generates unnatural body movements based on the given audio, as shown in our supplementary video, but the results of these metrics can not show this gap. Besides, the SRGR can not show the difference of various methods, because SRGR only considers the ground truth semantic motion for a specific speech content. All the same, the proposed method can outperform previous works across most metrics. Although we only consider two modalities (audio and speaker identity), the proposed method can excels state-of-the-art model that employs all five modalities, such as DiffStyleGesture [44].
Table 4: The statistical data of quantitative results by comparing SIGGesture to the baselines. The $\downarrow$" means the lower, the better. The $\uparrow$" means the higher, the better.
[…]
DiffuseStyleGesture (IJCAI’23) [44] 10.14 0.233 0.504 11.975
TalkSHOW (CVPR’23)[48] 7.313 0.279 0.463 12.859
QPGesture (CVPR’23)[45] 19.921 0.209 0.453 9.438
EMAGE (24) [31] 5.430 0.272 0.679 13.075
SIGGesture 2.021 0.263 0.707 14.020
4.6 Ablation study
In this subsection, we evaluate the contribution of two main parts, large-scale pre-training and semantic injection, by user study. The comparison results are shown in Table 5. The findings reveal that the pre-training with noise data can improve the quality of basic rhythmic gesture. In addition to pre-training, the proposed semantic injection can significantly improve the quality of semantic gesture.
Table 5: The statistical results of ablation study. "w/o" means "without".
[…]
baseline 0.56 0.63 0.50 0.38
w/o pre-training 0.64 0.72 0.72 0.87
w/o semantic injection 0.81 0.85 0.76 0.53
SIGGesture (full) 0.82 0.87 0.83 0.92
4.7 Generalization
To evaluate the generalizability, we test the proposed method on a broader set of audio samples, including out-of-domain English, Chinese and Japanese. The results are shown in Figure 5. It is also recommended to view the supplementary for more intuitive understanding and more examples. As shown in the visualization, the proposed method can accurately capture the key semantics and present precise gestures cross different language. The enhanced generalization capacity of our methodology is attributable to two principal components. Firstly, the employment of pseudo-labeled data can significantly improve the generalizability of rhythmic gestures. Secondly, disentangling the generation of rhythmic and semantic gestures by semantic injection can strengthen the controllability of semantic gesture synthesis. We have established an extensive collection of semantic gestures, encompassing a comprehensive range of gestural expressions. We leverage LLMs as a bridge to link the semantic gesture database and linguistic contexts, which has capable of robust generalization across diverse language. This method effectively addresses the challenges posed by the sparsity and long-tail distribution of semantic gestures. As a result, the proposed method has distinguished generalizability over different speech inputs.
Figure 5: The visual results of the proposed method on in-the-wild speeches (English, Chinese, and Japanese). The gestures in the red box are semantic gesture, which is labeled by the content in green box.
5 Conclusion
In this paper, we present a novel method named SIGGesture for high-quality co-speech gesture synthesis. To achieve this, we propose a semantic injection method with LLMs to improve controllability, interpretability, and generalization ability. With the ability of LLMs, this method can be applied to other language without any adjustment. Besides, we propose a large-scale dataset from Internet videos for pre-training the diffusion-based synthesis model. With the usage of pre-training and fine-tuning paradigm, the generated gestures are more robust to the variation of the audio. The proposed datasets and related statistical experiments can benefit the research of co-speech gesture synthesis, including controllable gesture synthesis, gesture foundation model in the future.

### v2

@@
 SIGGesture: Generalized Co-Speech Gesture Synthesis via Semantic Injection with Large-Scale Pre-Training Diffusion Models
-Abstract
-The automated synthesis of high-quality 3D gestures from speech is of significant value in virtual humans and gaming.
-Previous methods focus on synthesizing gestures that are synchronized with speech rhythm, yet they frequently overlook the inclusion of semantic gestures.
-These are sparse and follow a long-tailed distribution across the gesture sequence, making them difficult to learn in an end-to-end manner.
-Moreover, generating gestures, rhythmically aligned with speech, faces a significant issue that cannot be generalized to in-the-wild speeches.
-To address these issues, we introduce SIGGesture, a novel diffusion-based approach for synthesizing realistic gestures that are of both high quality and semantically pertinent.
-Specifically, we firstly build a strong diffusion-based foundation model for rhythmical gesture synthesis by pre-training it on a collected large-scale dataset with pseudo labels.
-Secondly, we leverage the powerful generalization capabilities of Large Language Models (LLMs) to generate proper semantic gestures for the various speech content.
-Finally, we propose a semantic injection module to infuse semantic information into the synthesized results during diffusion reverse process.
-Extensive experiments demonstrate that the proposed SIGGesture significantly outperforms existing baselines and shows excellent generalization and controllability.
-Keywords:
-Co-speech gesture synthesis Semantic gestures Large Language Models Diffusion models
-1 Introduction
-Body gesture can greatly improve people’s expressiveness in communication, which can emphasize or convey specific semantics with non-verb body movements [8, 18, 49, 36].
-Besides, high-quality 3D gesture animation can present vivid virtual characters that help listeners to concentrate and improve the intimacy between humans and virtual characters.
-Therefore, there is a growing need for automatic high-quality co-speech gesture synthesis in many applications, including Virtual Humans, non-player gaming characters, and robot assistants.
-Many researchers have proposed lots of approaches including rule-based methods [24, 35] and data-driven methods [52, 53, 2].
-Figure 1: The framework of the proposed method.
-The upper shows the training details of diffusion process and denoising process.
-The lower shows the inference process of synthesizing the gesture from the conditions.
-Specifically, given audio feature, speaker identity, and speech content, the audio feature and speaker identity are directly fed into the denoising network.
-The speech content are as input in the interaction with LLM, and the LLM will generate a set of candidate semantic gestures from the input text for the next semantic injection process.
-3D co-speech gesture generation aims to generate gestures with natural movements, accurate rhythm, and clear semantics based on a given speech.
-Rule-based methods [24, 35] achieve this goal by hard-rules, such as keyword "good" with a simple gesture "thump up".
-However, these approaches can only deal with simple semantic gestures, which are lack of naturalness, smoothness, and diversity.
-The principal reason for this is that generating co-speech gestures poses a challenge due to its inherently complex many-to-many mapping.
-Recently, deep learning with large-scale dataset has achieved many breakthroughs, especially in content synthesis.
-With the aid of deep learning, co-speech gesture synthesis also bloomed many remarkable methods.
-These methods attempt to solve this problem by training a conditional model with multi-modal inputs (audio, text, speaker identities).
-Co-speech gestures include rhythmic gesture and semantic gesture.
-Some works attempt to generate semantic gestures in co-speech gesture generation, such as GestureDiffuCLIP [2], LivelySpeaker [52], and Bodyformer [38].
-Both GestureDiffuCLIP [2] and LivelySpeaker [52] use CLIP [40] to learn the relationship between semantic text and motion.
-However, semantic gesture constitutes only a minor portion of the entire sequence and follows a long-tailed distribution, complicating the direct learning of its multi-modal relationships.
-Moreover, synthesizing gestures that are rhythmically synchronized with the given speech faces a challenge of generalizing to in-the-wild speech scenarios.
-Table 1: Comparison of Co-Speech Gesture Datasets. We ignore some datasets with 2D and 3D key-points label. "3D Rot" indicates 3D skeleton with bone rotation. "PGT" means Pseudo Ground Truth.
+Abstract.
+The automated synthesis of high-quality 3D gestures from speech holds significant value for virtual humans and gaming.
+Previous methods primarily focus on synchronizing gestures with speech rhythm, often neglecting semantic gestures.
+These semantic gestures are sparse and follow a long-tailed distribution across the gesture sequence, making them challenging to learn in an end-to-end manner.
+Additionally, generating rhythmically aligned gestures that generalize well to in-the-wild speech remains a significant challenge.
+To address these issues, we introduce SIGGesture, a novel diffusion-based approach for synthesizing realistic gestures that are both high-quality and semantically pertinent.
+Specifically, we firstly build a robust diffusion-based foundation model for rhythmical gesture synthesis by pre-training it on a collected large-scale dataset with pseudo labels.
+Secondly, we leverage the powerful generalization capabilities of Large Language Models (LLMs) to generate appropriate semantic gestures for various speech transcripts.
+Finally, we propose a semantic injection module to infuse semantic information into the synthesized results during the diffusion reverse process.
+Extensive experiments demonstrate that SIGGesture significantly outperforms existing baselines, exhibiting excellent generalization and controllability.
+Co-speech gesture synthesis, Semantic gestures, Large Language Models, Diffusion models
+journalyear: 2024copyright: acmlicensedconference: SIGGRAPH Asia 2024 Conference Papers; December 3–6, 2024; Tokyo, Japanbooktitle: SIGGRAPH Asia 2024 Conference Papers (SA Conference Papers ’24), December 3–6, 2024, Tokyo, Japandoi: 10.1145/3680528.3687677isbn: 979-8-4007-1131-2/24/12ccs: Computing methodologies Animationccs: Computing methodologies Natural language processingccs: Computing methodologies Neural networks
+1. Introduction
+Body gesture can significantly enhance people’s expressiveness in communication by emphasizing specific intentions or conveying specific semantics (Cassell et al., 1999; Goldin-Meadow and Alibali, 2013; Yoon et al., 2020; Nyatsanga et al., 2023).
+Additionally, high-quality 3D gesture animation can present vivid virtual characters that help listeners to concentrate and improve the intimacy between humans and virtual characters.
+Consequently, there is a growing demand for automatic high-quality co-speech gesture synthesis in various applications, including virtual humans, non-player gaming characters, and robot assistants.
+Many researchers have proposed lots of approaches including rule-based methods (Huang and Mutlu, 2012; Marsella et al., 2013) and data-driven methods (Zhi et al., 2023; Zhu et al., 2023; Ao et al., 2023).
+Figure 1.
+The framework of the proposed method is illustrated in two parts.
+The upper section details the training processes of diffusion and denoising.
+The lower part demonstrates the inference process for synthesizing gestures based on the given conditions.
+Specifically, the audio features and speaker identity are directly fed into the denoising network.
+Meanwhile, the speech textual content are used as input in the interaction with the LLM, which generates a set of candidate semantic gestures for the subsequent semantic injection process.
+3D co-speech gesture generation aims to produce gestures with natural movements, accurate rhythm, and clear semantics based on the given speech.
+Rule-based methods (Huang and Mutlu, 2012; Marsella et al., 2013) achieve this by using predefined rules, such as associating the keyword ”good” with a simple ”thumbs up” gesture.
+However, these approaches, lacking naturalness, smoothness, and diversity, can only handle simple semantic gestures.
+The main reason is that generating co-speech gestures is challenging due to the inherently complex many-to-many mapping.
+Recently, deep learning with large-scale datasets has achieved significant breakthroughs, especially in content synthesis.
+With the aid of deep learning, co-speech gesture synthesis has seen the emergence of many remarkable methods.
+These methods attempt to solve this problem by training a conditional model with multi-modal inputs (audio, text, speaker identities, etc.).
+Co-speech gestures include rhythmic gestures and semantic gestures.
+Some works attempt to generate high-quality co-speech gesture that contains accurate semantic gestures , such as GestureDiffuCLIP (Ao et al., 2023), LivelySpeaker (Zhi et al., 2023), and Bodyformer (Pang et al., 2023).
+Both GestureDiffuCLIP (Ao et al., 2023) and LivelySpeaker (Zhi et al., 2023) use CLIP (Radford et al., 2021) to learn the relationship between semantic transcripts and motion.
+However, semantic gestures constitute only a minor portion of the entire sequence and follow a long-tailed distribution, complicating the direct learning of their multi-modal relationships.
+Moreover, synthesizing gestures rhythmically synchronized with speech faces the challenge of generalizing to in-the-wild scenarios.
+Table 1. Comparison of co-speech gesture datasets. We ignore some datasets with 2D and 3D key-points label. ”3D Rot” indicates 3D skeleton with bone rotation. ”PGT” means Pseudo Ground Truth.
 Trinity
-Mocap
-2017 [14]
+Mocap 2017
+(Ferstl and McDonnell, 2018)
 TWH
-Mocap
-2019 [28]
+Mocap 019
+(Lee et al., 2019)
 Hahibie et al.
-S2G-3D
-2021 [20]
+S2G-3D 2021
+(Habibie et al., 2021)
 BEAT(X)
-Mocap
-2022 [32]
+Mocap 2022
+(Liu et al., 2022b)
 ZEEG
-Mocap
-2022 [16]
+Mocap 2022
+(Ghorbani et al., 2023)
 Yoon. et al.
-TED-SMPL
-2023 [34]
+TED-SMPL 2023
+(Lu et al., 2023)
 Talkshow
-S2G-SMPL
-2023 [48]
-LSMoG
+S2G-SMPL 2023
+(Yi et al., 2023)
+Gesture400
 Ours
@@
 Duration (hours) 4 20 38 76 4 30 27 393
-For addressing the after-mentioned issues, we propose a diffusion-based semantic injection method with large language model for semantic gesture synthesis, which has excellent controllability, interpretability, and generalization ability.
-It begins with a insight that real-word human conversation contains limited semantic gestures.
-Besides, for a specific semantic word, such as "wonderful", different people usually present similar gesture such as "opening both arms".
-During the diffusion reverse process, we can utilize semantic matched gestures as prompts to guide the generative model in synthesizing the appropriate semantic gestures.
-Therefore, we collect a large-scale semantic motion database, which can include most semantic motion in communication.
-Large Language Models (LLMs), such as GPT3 [6] and GPT-4, have shown their ability not only as proficient models for natural language processing (NLP) tasks but also as potent instruments for tackling complex problems.
-Therefore, we introduce large-language model to assign contextually appropriate gestures to the corresponding text.
+To address the aforementioned issues, we propose a diffusion-based semantic injection method with large language model for semantic gesture synthesis, which offers excellent controllability, interpretability, and generalization.
+It begins with the insight that real-world human conversation contains relatively limited semantic gestures.
+For example, for specific semantic words like ”wonderful,” different people usually present similar gestures, such as ”opening both arms.” During the diffusion reverse process, we can utilize semantically matched gestures as prompts to guide the generative model in synthesizing appropriate semantic gestures.
+Therefore, we collect a large-scale semantic motion database that includes most semantic motions in communication.
+LLMs, such as GPT3 (Brown et al., 2020) and GPT-4 (Achiam et al., 2023), have demonstrated their ability not only as proficient models for natural language processing (NLP) tasks but also as potent instruments for tackling complex problems.
+Therefore, we introduce large-language models to assign contextually appropriate gestures to the corresponding text.
 Leveraging the strong semantic analysis capabilities of LLMs, the proposed method is adept at handling various languages.
-For a specific speech, different people will present notably different rhythmic gesture.
-For such many-to-many issue, the best solution is to use deep learning with large-scale dataset to learn their complex relationship.
-Therefore, it is necessary to build a robust foundational model for rhythmic gesture.
-Then, the semantic gesture can be injected into the rhythmic gesture with naturalness.
-However, the co-speech gesture generation research community lack of 3D skeletal data because of expensive motion capture system, which is different from other synthesis tasks, such as text2image, text2speech.
-Therefore, for improving the basic quality of rhythmic gesture, we collect a large-scale gesture dataset named LSMoG for pre-training that reaches to 400 hours.
-Table 1 shows the statistical data of co-speech gesture datasets.
-Some datasets with 2D and 3D key-points label are ignored because they are difficult to use in real applications.
-To the best of our knowledge, it is the largest dataset in the research of co-speech gesture synthesis field.
-By the optimization of pre-traing and fine-tuning, the synthesized gesture motion has excellent naturalness and generalization.
-The main contributions of our works contains the following aspects:
-  1. 1.
-    We propose a LLMs-based semantic augmentation method for co-speech gesture synthesis, which can greatly improve the generalization ability in semantic gesture synthesis.
-For controllable semantic gesture synthesis, we build a set of semantic gesture clips and use LLMs to generate proper candidate gestures for a specific speech content.
-  2. 2.
-    We build a strong diffusion-based foundation model for rhythmical gesture synthesis by pre-training it on a collected large-scale dataset with pseudo labels.
-This dataset is the largest dataset for co-speech gesture synthesis, which contains approximately 400 hours of motion sequences.
-  3. 3.
+For a specific speech, different people will present notably different rhythmic gestures.
+For such many-to-many issue, the best solution is to use deep learning with a large-scale dataset to learn their complex relationships.
+Therefore, it is necessary to build a robust foundational model that can synthesize natural and smooth rhythmic gestures.
+Then, the semantic gestures can be naturally injected into the rhythmic gestures.
+However, the co-speech gesture generation research community lacks 3D skeletal data due to the expensive motion capture system, which is different from other synthesis tasks, such as text-to-image, text-to-speech.
+Therefore, for improving the basic quality of rhythmic gestures, we collected a large-scale gesture dataset named Gesture400 for pre-training, which contains approximately 400 hours of data.
+Table 1 shows the statistical data of these co-speech gesture datasets.
+Some datasets with 2D and 3D key-point labels are ignored because they are difficult to use in real applications.
+To the best of our knowledge, the proposed dataset is the largest dataset in the field of co-speech gesture synthesis research.
+By the optimization of pre-training and fine-tuning, the synthesized gesture motion exhibits excellent naturalness and generalization.
+The main contributions of our work contain the following aspects:
+  1. (1)
+    We propose an LLMs-based semantic augmentation method for co-speech gesture synthesis, which can greatly improve the generalization ability in semantic gesture synthesis.
+For controllable semantic gesture synthesis, we build a set of semantic gesture clips and use LLMs to generate appropriate candidate gestures for the specific speech transcripts.
+  2. (2)
+    We build a robust diffusion-based foundational model for rhythmical gesture synthesis by pre-training it on a collected large-scale dataset with pseudo labels.
+This dataset is the largest dataset for co-speech gesture synthesis, containing approximately 400 hours of motion sequences.
+  3. (3)
     Extensive experiments show that the proposed method outperforms state-of-the-art methods by a large margin.
-Especially, the visualization comparison indicates that our method has more stable, expressive, and robust results than other approaches.
-2 Related Work
-2.1 Co-Speech Gesture Generation
-The research of co-speech gesture generation can be divided into two branches, Rule-based methods [9, 35, 10] and learning-based methods [1, 2, 33, 17, 33].
-Rule-based works [9, 35, 10] for co-speech gesture synthesis are mainly generated from a pre-defined motion database by keyword matching or other specific rules.
-These rule-based based methods require lots of human effort in defining gesture units and complex mapping rules, which is costly and inefficient.
+In particular, the visualization comparisons indicate that our method produces more stable, expressive, and robust results than other approaches.
+2. Related Work
+2.1. Co-Speech Gesture Generation
+The research of co-speech gesture generation can be divided into two branches, rule-based methods (Cassell et al., 1994; Marsella et al., 2013; Cassell et al., 2001) and learning-based methods (Ao et al., 2022, 2023; Liu et al., 2022a; Ginosar et al., 2019).
+Rule-based works (Cassell et al., 1994; Marsella et al., 2013; Cassell et al., 2001) for co-speech gesture synthesis are mainly generated from a pre-defined motion database by keyword matching or other specific rules.
+These rule-based methods require lots of human effort in defining gesture units and complex mapping rules, which are costly and inefficient.
 Besides, the results of rule-based methods are usually lack of smoothness and naturalness.
 Co-speech gesture synthesis is a complex problem that requires consideration of the audio, the gesture, and their many-to-many relationships.
-Benefit from recent advanced developments of deep learning, all of recent methods [1, 2, 33] adopt deep neural networks as a strong tool to learn the complex relationships from audio to gesture in an end-to-end manner.
-Earlier works [17, 39] explore various network architecture to regress 2D keypoints of the human body movements because of lacking high-quality 3D gesture datasets.
-These 2D datasets are constructed by using off-the-shelf pose estimator [7] to obtain pseudo 2D gesture annotations from the online videos.
-However, these 2d results are lack of expressiveness in real application.
-Recently, some high-quality speech-gesture corpus such as BEAT [32] and ZeroEGGS [16] are released, which contribute to the development of co-speech gesture synthesis.
-Existing works focus on exploring the effectiveness of different network architectures, including the multi-layer perceptron (MLP) [27], convolutional neural networks (CNN) [20, 48], recurrent neural networks (RNNs) [49, 22, 33, 3] and Transformers [4, 38].
-Besides the exploration of model architecture, some approaches [17, 44] dive into researching the connections between co-speech gesture and speech audio, text transcript, speaking style, and speaker identity.
-Besides, some approaches involve the adversarial training [17, 33], Phase-Guided Motion Matching[45], and reinforcement learning [42] to guarantee realistic results.
-In addition to gestures, some methods [48, 11, 31] begin to explore the generation of full-body movements, including body gestures and facial expressions.
-The rapid development of diffusion models [23, 41] has surpassed the traditional synthesis paradigm such as GAN [19] and VAE [26], showcasing a powerful and realistic characteristic in synthesis.
-Diffusion-based motion generation also become a popular research direction, such as text2motion [50].
+Benefit from recent advanced developments of deep learning, all of recent methods (Ao et al., 2022, 2023; Liu et al., 2022a) adopt deep neural networks as a strong tool to learn the complex relationships from audio to gesture in an end-to-end manner.
+Earlier works (Ginosar et al., 2019; Qian et al., 2021) explore various network architecture to regress 2D keypoints of the human body movements because of lacking high-quality 3D gesture datasets.
+These 2D datasets are constructed by using off-the-shelf pose estimator (Cao et al., 2017) to obtain pseudo 2D gesture annotations from the online videos.
+However, The 2D results face significant challenges in practical applications..
+Recently, some high-quality speech-gesture corpus such as BEAT (Liu et al., 2022b) and ZeroEGGS (Ghorbani et al., 2023) are released, which contribute to the development of co-speech gesture synthesis.
+Some of existing works focus on exploring the effectiveness of different network architectures, including the Multi-Layer Perceptron (MLP) (Kucherenko et al., 2020), Convolutional Neural Networks (CNN) (Habibie et al., 2021; Yi et al., 2023), Recurrent Neural Networks (RNNs) (Yoon et al., 2020; Hasegawa et al., 2018; Liu et al., 2022a; Bhattacharya et al., 2021a) and Transformers (Bhattacharya et al., 2021b; Pang et al., 2023).
+Besides the exploration of model architecture, some approaches (Ginosar et al., 2019; Yang et al., 2023b) dive into researching the connections between co-speech gesture and speech audio, text transcript, speaking style, and speaker identity.
+Furthermore, some approaches involve the adversarial training (Ginosar et al., 2019; Liu et al., 2022a), phase-guided motion matching(Yang et al., 2023a), and reinforcement learning (Sun et al., 2023) to guarantee realistic results.
+In addition to gestures, some methods (Yi et al., 2023; Chen et al., 2024; Liu et al., 2023) begin to explore the generation of full-body movements, including body gestures and facial expressions.
+The rapid development of diffusion models (Ho et al., 2020; Song et al., 2020) has surpassed the traditional synthesis paradigm such as GAN (Goodfellow et al., 2020) and VAE (Kingma and Welling, 2013), showcasing a powerful and realistic characteristic in synthesis.
+Diffusion-based motion generation also become a popular research direction, such as text2motion (Zhang et al., 2024b).
 To be specific, text-guided motion can synthesize realistic and expressive results with diffusion model.
-In co-speech gesture synthesis, recent approaches such as GestureDiffuCLIP[2], LivelySpeaker [52], DiffStyleGesture [44], C2G2[25] are diffusion-based approaches.
-2.2 Semantic-aware Co-Speech Gesture Generation
+In co-speech gesture synthesis, recent approaches such as GestureDiffuCLIP(Ao et al., 2023), LivelySpeaker (Zhi et al., 2023), DiffStyleGesture (Yang et al., 2023b), C2G2(Ji et al., 2023), LDA (Alexanderson et al., 2023), Remodiffuse (Zhang et al., 2023a) are diffusion-based approaches.
+2.2. Semantic-aware Co-Speech Gesture Generation
 In gesture language, semantic gesture units are crucial for conveying specific intentions, thoughts, emotions, and enhancing gestural expressiveness.
-For generating more expressive gestures, there are also some attempts such as Gesticulator [27], GestureDiffuCLIP[2], LivelySpeaker[52], Bodyformer [38].
-Specifically, GestureDiffuCLIP[2] uses CLIP [21] to align motion and text semantic, then use the text as additional condition to control the semantic gestures.
-LivelySpeaker[52] is a two-stage framework for semantic and rhythmic gesture generation, which aim at decoupling the semantic gesture generation and rhythmic gesture generation.
+For generating more expressive gestures, there are also some attempts such as Gesticulator (Kucherenko et al., 2020), GestureDiffuCLIP(Ao et al., 2023), LivelySpeaker(Zhi et al., 2023), Bodyformer (Pang et al., 2023).
+Specifically, GestureDiffuCLIP(Ao et al., 2023) adopts CLIP (Hafner et al., 2021) to align motion and text semantic, then use the textual feature as additional condition to control the generation process.
+LivelySpeaker(Zhi et al., 2023) is a two-stage framework for semantic and rhythmic gesture generation, which aim at decoupling the semantic gesture generation and rhythmic gesture generation.
 In semantic gesture generation, they also use a CLIP style module to align the semantic motion and semantic text.
-Some approaches add semantic text as additional input to fuse semantic into the model, such as Rhythmic Gesticulator [1], Gesticulator [27], FreeTalker [46] and Trimodal [49].
-BodyFormer [38] takes both low-level and high-level speech features as input and generates a sequence of realistic 3D body gestures in an auto-regressive manner.
+Besides, some approaches add semantic text as additional input to fuse semantic into the model, such as Rhythmic Gesticulator (Ao et al., 2022), Gesticulator (Kucherenko et al., 2020), FreeTalker (Yang et al., 2024) and Trimodal (Yoon et al., 2020).
+BodyFormer (Pang et al., 2023) takes both low-level and high-level speech features as input and generates a sequence of realistic 3D body gestures in an auto-regressive manner.
 The high-level speech features represent the semantic gestures.
-With the development of LLMs, some approaches also use it to extract motion-related information from textual input, such as GesGPT[15], MoConVQ [47].
-3 Method
-3.1 Problem Formulation
+With the development of LLMs, some approaches also use it to extract motion-related information from textual input, such as GesGPT(Gao et al., 2023), MoConVQ (Yao et al., 2023).
+The proposed method uses Large Language Models (LLMs) to generate proper mapping between the text transcripts and semantic gesture units, which shows excellent performance in both generalization and controllability.
+In semantic gesture generation, Semantic Gesticulator(Zhang et al., 2024a) is the most similar work to ours in the same period.
+Semantic Gesticulator(Zhang et al., 2024a) also attempts to synthesize semantically accurate gestures by introducing a LLM-based semantic gesture units retrieval from a predefined dataset.
+The gesture generator of Semantic Gesticulator(Zhang et al., 2024a) is based on the GPT-2 (Radford et al., 2019), which predicts the future gesture tokens in an auto-regressive manner.
+Compared with their work, we use a strong diffusion-based method as fundamental gesture generator to fuse semantic gestures and rhythmic gestures.
+3. Method
+3.1. Problem Formulation
 Given a specific audio, its corresponding text, and identity information, the proposed method can generate high-quality 3D skeletal co-speech gesture results.
-To be specific, considering a sample of raw audio ${A=\{a_{0},a_{1},...,a_{N}\}}$, we adopt Jukebox [13] to extract its acoustic feature, which is trained on large-scale music datasets.
-All of the motion sequences =$\mathcal{M}=\{m_{0},m_{1},...,m_{T}\}$ are encoded into discrete tokens by VQVAE[37].
-With VQVAE, a motion can be represented by latent embedding $\hat{e}=\{e_{0},e_{1},...,e_{T}\}$.
-Then, we adopt transformer-based diffusion model as the denoising network to recover the latent embedding by removing the noise.
-The reverse denoising process of the optimized diffusion model $G$ parameterized by $\theta$ can recover the motion latent embedding $\hat{e}=\{e_{0},e_{1},...,e_{T}\}$ conditioned on the speech audio sequence and identity information.
-With integration with LLM, it can provide accurate candidate semantic gestures for the speech content.
-Finally, the candidate semantic gesture embedding are fused into the synthesized motion latent embedding by semantic injection module.
-3.2 Semantic Gesture Collection
+To be specific, considering a sample of raw audio $A$, we adopt Jukebox (Dhariwal et al., 2020) to extract its acoustic feature, which is pre-trained on large-scale music datasets.
+All of the motion sequences $\mathcal{M}$ are encoded into discrete tokens by VQVAE (Oord et al., 2017).
+Details of VQVAE can be found in the appendix.
+By VQVAE encoding, a motion can be represented by a latent embedding $x_{0}$ Then, we adopt transformer-based diffusion model as the denoising network to recover the latent embedding by removing the noise.
+The reverse denoising process of the optimized diffusion model $G$ parameterized by $\theta$ can recover the motion latent embedding $x_{0}$ conditioned on the speech audio sequence and identity information.
+With integration with LLM, it can provide accurate candidate semantic gestures for the speech transcripts.
+Then, the candidate semantic gesture embedding are fused into the synthesized motion latent embedding by semantic injection module.
+Finally, the synthesized motion are decoded to continues 3D skeleton rotation by the pre-trained VQVAE.
+3.2. Semantic Gesture Collection
+Figure 2. The statistical data of semantic gesture dataset (upper left part), some examples of semantic label (lower left part), and some semantic gesture examples (right part).
 Through extensive observation of numerous instances, it has been discerned that semantic gestures maintain a universality in gestural communication.
-Specifically, it is evident that different individuals display similar gestural expressions when articulating identical semantic concepts.
+Specifically, it is evident that different individuals display similar gestural expressions when articulating identical semantic concepts (Lascarides and Stone, 2009; Zhang et al., 2024a).
 This observation provides a feasible methodology for the collection of an extensive array of semantic gesture actions, aimed at covering a comprehensive range of semantic contexts.
-We collect a large-scale semantic motion clip $G=\{G_{0},G_{1},...,G_{N}\}$ and text $T=\{T_{0},T_{1},...,T_{N}\}$ pairs from BEAT dataset by manual selection.
-$N$ is the number of the semantic gestures.
-All of the semantic clip is shorter than 3 seconds and each motion clip represent a whole semantic motion.
-After selection, we have collected 2,537 semantic motion clips.
-Some examples of semantic gestures are shown in Figure 2.
+We collect a large-scale semantic gesture dataset that contains motion clips $G=\{G_{0},G_{1},...,G_{N}\}$ and corresponding text $T=\{T_{0},T_{1},...,T_{N}\}$ by manual selection and modification with the aid of technical artist.
+Some of these semantic gestures are borrowed from BEAT dataset.
+All of the semantic motion clips are shorter than 3 seconds and each motion clip is a whole independent semantic motion.
+After manual selection and check, we have collected 2,537 semantic motion clips.
+Some examples of semantic gestures and the distribution statistics are shown in Figure 2.
+More examples can be found in the supplementary video.
 These motion clips are sorted into a predefined motion semantic system.
-Figure 2: The statistical data of semantic gesture dataset (upper left part), some examples of semantic label (lower left part), and some semantic gesture examples (right part).
-All of semantic gesture clips are encoded into latent embedding $V=\{v_{0},v_{1},...,v_{N}\}$ by a pre-trained VQVAE model.
+All of semantic gesture clips are encoded into latent embedding $V_{semantic}=\{v_{0},v_{1},...,v_{N}\}$ by a pre-trained VQVAE model.
 To improve the diversity of semantic gesture, we add a random noise to the embeddings during semantic injection.
-3.3 Diffusion models with semantic injection
+3.3. Diffusion models with semantic injection
 SIGGesture employs a diffusion model to synthesize the latent representation of gesture, which consists of a diffusion process and a denoising process, as shown in upper part of Figure 1.
-Given a specific real motion data distribution $p(x_{0})$, our goal is learn a distribution $q_{\theta}(x_{0})$ that is parameterized by ${\theta}$.
-Diffusion process.
-We follow the previous DDPM [23] definition of diffusion as a Markov noising process with latents $x_{0:T}$ that follow a forward noising process $p(x_{0:T}|x_{0})$.
-Here, $x_{0}$ is the latent motion data and same as previous notation $\hat{e}$.
-The forward diffusion process is defined as a Markov chain that gradually adds Gaussian noise to the latent representation of motion, as follows
-$\begin{split}q(x_{0:T}|x_{0})=\prod_{t=1}^{T}q(x_{t}|x_{t-1}),\end{split}$ (1)
+Given a specific real motion data distribution $p(x_{0})$, our goal is learn a distribution $q_{\theta}(x_{0})$ that is parameterized by ${\theta}$ to approximate the real distribution.
+3.3.1. Diffusion process.
+We follow the previous DDPM (Ho et al., 2020) definition of diffusion as a Markov noising process with latent $x_{0:T}$ that follows a forward noising process $p(x_{0:T}|x_{0})$.
+Here, $x_{0}$ is the latent embedding of real motion data.
+The forward diffusion process is defined as a Markov chain that gradually adds Gaussian noise to the latent representation of real motion, as follows
+(1) $\begin{split}q(x_{0:T}|x_{0})=\prod_{t=1}^{T}q(x_{t}|x_{t-1}),\end{split}$
 where
-$\begin{split}q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}{I}).\end{split}$ (2)
-$\beta_{t}$ are variance schedule hyperparameters that controls the distribution of Gaussian noise.
+(2) $\begin{split}q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}{I}).\end{split}$
+$\beta_{t}$ are variance schedule hyperparameters that control the distribution of Gaussian noise.
 In the reverse denoising process, $\beta_{t}$ are pre-defined constant parameters.
-By this diffusion process, the original latent representation of motion $x_{0}$ progressively substituted by random noises.
+In the diffusion process, the original latent representation of motion $x_{0}$ progressively substituted by random noises.
 When $T$ approaches infinity, the distribution of $x_{T}$ follows a pure white noise distribution.
-Denoising process.
+3.3.2. Denoising process.
 Through diffusion process, the real data are transferred into pure white noise by adding random noise.
 On the contrary, the denoising process recovers the real data from pure white noise by removing the random noise.
-The unconditional reverse process of recover a gesture motion can be represented by the following formula
-$\begin{split}p_{\theta}(x_{0:T})=p_{\theta}(x_{0})\prod_{t=1}^{T}p_{\theta}(x_{t-1}|x_{t}),\end{split}$ (3)
+The unconditional reverse process of recover a gesture motion embedding can be represented by the following formula
+(3) $\begin{split}p_{\theta}(x_{0:T})=p_{\theta}(x_{0})\prod_{t=1}^{T}p_{\theta}(x_{t-1}|x_{t}),\end{split}$
 where
-$\begin{split}p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t};u_{\theta}(x_{t},t),\beta_{t}I).\end{split}$ (4)
-The noise $x_{t}$ is sampled from the following defined distribution
-$\begin{split}p(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha_{t}}},(1-\bar{\alpha_{t}}){I}).\end{split}$ (5)
+(4) $\begin{split}p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t};u_{\theta}(x_{t},t),\beta_{t}I).\end{split}$
+The intermediate state $x_{t}$ is sampled from the following defined distribution
+(5) $\begin{split}p(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha_{t}}},(1-\bar{\alpha_{t}}){I}).\end{split}$
 where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha_{t}}=\prod_{k=0}^{t}\alpha_{k}$.
@@
 The details of the denoising network are shown in the upper right part of Figure 1.
-In our method, the conditions are audio $c_{audio}$ and speaker identity $c_{id}$ respectively.
+In our method, the conditions contains audio $c_{audio}$ and speaker identity $c_{id}$.
 For simple notation, we mark the two conditions as $c=[c_{audio},c_{id}]$.
-It should be noted that the audio feature is extracted by Jukebox [13] instead of WavLM [12].
-Jukebox [13], trained on large-scale music datasets, can focus on the rhythmic features rather than the speech semantic content.
-By injecting the condition $c$ into the generation process, the reverse process of each time step can be updated by the following formula
-$\begin{split}p_{\theta}(x_{t-1}|x_{t},c)=\mathcal{N}(x_{t};u_{\theta}(x_{t},t,c),\beta_{t}I).\end{split}$ (6)
+It should be noted that the audio feature is extracted by Jukebox (Dhariwal et al., 2020) instead of WavLM (Chen et al., 2022).
+Jukebox (Dhariwal et al., 2020), trained on large-scale music datasets, can focus on the rhythmic features rather than the speech semantic transcripts.
+By injecting the condition $c$ into generation, the reverse process of each time step $t$ can be updated by the following formula
+(6) $\begin{split}p_{\theta}(x_{t-1}|x_{t},c)=\mathcal{N}(x_{t};u_{\theta}(x_{t},t,c),\beta_{t}I).\end{split}$
 In a short summary, we firstly start denosing process by sampling $x_{T}$ from a pure white nose distribution $\mathcal{N}(0,I)$.
 Then, we iteratively denoise the latent variable $x_{t}$ to obtain the final results $x_{0}$.
-Training objective.
+3.3.3. Training objective.
 For a speech condition $c$, we reverse the forward diffusion process by learning to estimate $p(x_{t},t,c)$ approximate $x_{0}$ with a parameterized model for all times step $t$.
-We use a loss function proposed in [23] to optimize the whole model, as following,
-$\begin{split}\mathcal{L}={E_{x,t}[\parallel x-p(x_{t},t,c)\parallel^{2}_{2}]}.\end{split}$ (7)
+We use a loss function proposed in (Ho et al., 2020) to optimize the whole model, as following,
+(7) $\begin{split}\mathcal{L}={E_{x,t}[\parallel x-p(x_{t},t,c)\parallel^{2}_{2}]}.\end{split}$
 During training, the model is trained under conditional and unconditional manner with a specific probability.
-Once the training converges, the the diffusion model can predict the noise parameters by considering the unconditional setting and condition $c$.
-Inference with semantic injection.
+Once the training converges, the diffusion model can predict the noise parameters by considering the unconditional setting and condition $c$.
+3.3.4. Inference with semantic injection.
 The inference process involves solely the denoising process of diffusion models.
 The pipeline of inference process is shown in the lower part of Figure 1.
-To be specific, speech with its corresponding text will be feed into the generation pipeline.
-It should be noted that the words in the text is aligned with speech in timeline.
-Each words have a duration time.
-Then, the text are fed into the LLMs to produce candidate semantic gestures $\mathcal{G}=(G_{0},...,G_{t})$ for the speech content.
+To be specific, a speech with its corresponding text will be fed into the generation pipeline.
+It should be noted that the words in the text are aligned with speech in the timeline.
+Each word has a duration time.
+Then, the texts are fed into the LLMs to produce candidate semantic gestures $G_{candicate}=(g_{0},...,g_{N})$ with a pre-defined prompt.
+$N$ is the number of semantic gestures in this text.
 The interaction with LLMs process is shown in the left-top corner of inference pipeline in Figure 1.
-With the ability of LLM, the proposed method can process various language, such as Chinese and Japanese.
-The candidate semantic gesture is encoded into latent embedding $\mathcal{P}=(p_{0},...,p_{n})$, where $n$ is number of candidate semantic gestures.
+With the ability of LLMs, the proposed method can process various languages, such as Chinese and Japanese.
+The candidate semantic gestures have been encoded into latent embedding.
 We set a control $K$ to determine the degree of semantic injection.
 From time step $T$ to $K$, the denoising latent variable is marked as $\hat{x_{t}}$ that contains the semantic information.
-$\begin{split}p_{\theta}(x_{K:T})=p_{\theta}(x_{T})\prod_{t=K}^{T}p_{\theta}(x_{t-1}|\hat{x}_{t}).\end{split}$ (8)
-The semantic gesture $P$ can be injected into the latent variable $x_{t}$ by the follow formula
-$\begin{split}\hat{x}_{t}=x_{t}*m+P*(1-m),\end{split}$ (9)
+(8) $\begin{split}p_{\theta}(x_{K:T})=p_{\theta}(x_{T})\prod_{t=K}^{T}p_{\theta}(x_{t-1}|\hat{x}_{t}).\end{split}$
+The semantic gesture $G_{candidate}$ can be injected into the latent variable $x_{t}$ by the follow formula
+(9) $\begin{split}\hat{x}_{t}=x_{t}*m+G_{candidate}*(1-m),\end{split}$
 where $m$ denotes the timeline mask of semantic words.
-From time step $K$ to 1, the denoising latent variable $x_{t}$ is not with semantic injection, which can preserve the structure and smoothness of the generated gesture.
-Generating long sequence.
-For synthesizing long motion sequence, the common practice is autoregressive method, which generates content sequentially.
+From time step $K$ to 1, the denoising latent variable $x_{t}$ does not contain semantic injection step, which can preserve the structure and smoothness of the generated gesture and improve the diversity of the results.
+3.3.5. Generating long sequence.
+For synthesizing long motion sequence, the common practice uses auto-regressive manner, which generates content sequentially.
 However, generative models trained on small pieces may cause unnaturalness and defectiveness.
-Instead of auto-regressive manner, we adopt DiffCollage [51] to generate long motion sequence.
-3.4 Large-scale motion corpus collection
+Instead of auto-regressive manner, we adopt DiffCollage (Zhang et al., 2023b) to generate long motion sequence.
+DiffCollage is a scalable probabilistic model that can synthesize large content in parallel.
+Specifically, we denote a long sequence as ${\boldmath{m}=[m^{0},m^{1},m^{2}]}$, where $m^{2}$ is the out-painted sequence generated by the conditional model $m^{2}|m^{1}$.
+Notably, this procedure makes a conditional independence assumption that $q(m^{2}|m^{0},m^{1})=q(m^{2}|m^{1})$.
+(10) $\begin{split}q(m)=q(m^{0},m^{1},m^{2})=q(m^{0},m^{1})q(m^{2}|m^{1}).\end{split}$
+Further, we can obtain the following formula,
+(11) $\begin{split}q(m)=\frac{q(m^{2},m^{1})q(m^{1},m^{0})}{q(m^{1})}.\end{split}$
+The score function of $q(m)$ in diffusion model can be represented as a sum over the scores of shorter gesture motions.
+(12) $\begin{split}\nabla q(m)=\nabla q(m^{2},m^{1})+\nabla q(m^{1},m^{0})-\nabla q(m^{1}).\end{split}$
+From formula 12, it can be found that we can synthesize different gesture clips in parallel since all individual scores can be calculated independently.
+Then, then we can merge the scores to compute the score of the target long audio condition.
+For more technical details, please refer to DiffCollage (Zhang et al., 2023b).
+3.4. Large-Scale Motion Corpus Collection
 In 3D animation, high-quality data collection is remarkably expensive, professional, and time-consuming.
@@
 In recent years, several high-quality datasets are released to the research community, as shown in Table 1.
-BEAT [32] is the largest dataset for co-speech gesture synthesis, which contains 76 hours data.
+BEAT (Liu et al., 2022b) is the largest dataset for co-speech gesture synthesis, which contains 76 hours data.
 However, data-driven approaches with larger models are also short for data.
-In the inspiration of large-language training paradigm, noise data for pre-training and high-quality data for fine-tuning, we aim at collecting a large-scale gesture corpus for pre-training by video motion capture technique such as HybridIK [29], SMPLify [5].
-Before running the video motion capture approach, we collect a large mount of videos about speech or talk from the Internet, such as TED.
-Then, we conduct the data cleaning pipeline to obtain the final shots of interests.
+In the inspiration of large-language training paradigm, noise data for pre-training and high-quality data for fine-tuning, we aim at collecting a large-scale gesture corpus for pre-training by video motion capture technique such as HybridIK (Li et al., 2021a), SMPLify (Bogo et al., 2016).
+To this end, we collect a large mount of videos about speech or talk from TED  [^1] https://www.ted.com and YiXi  [^2] https://yixi.tv.
+Then, we conduct the data post-processing pipeline to obtain the final shots of interests.
 Table 2 shows the statistical data of the motion corpus collection.
@@
 As shown in Table 1, the proposed dataset is significantly larger than previous datasets by at least 6 times.
-Table 2: Statistical data of the motion corpus collection.
+Table 2. Statistical data of the collected motion corpus collection.
 Items Number Average length Total length
@@
 final results 70,784 $\sim$20s $\sim$393h
-4 Experiments
-4.1 Datasets
+Figure 3. Visualization results between the proposed method and other state-of-the-art methods.
+4. Experiments
+4.1. Datasets
 In this paper, we use the largest high-quality speech-gesture dataset BEAT to evaluate the proposed method.
@@
 It contains about 76 hours motion-audio paired data of 30 speakers talking about different topics.
-We deeply clean the dataset by removing some defective motion sequences such as speaker drinking, wandering, and incorrect motion.
-After cleaning the raw data, the BEAT dataset contains 46 hours motion data and its corresponding text and audio.
+We thoroughly post-process the dataset by removing some defective motion sequences such as speaker drinking, wandering, and incorrect motion.
+Then, the processed BEAT dataset contains 46 hours motion data and its corresponding text and audio.
 During training the VQVAE model, the original Euler angles are converted to rotation matrix for better convergence.
 During training the diffusion-based generation model, all motions are represented by latent embeddings in VQVAE latent space.
-The datasets collected from Internet by video motion capture are only used for pre-training because of its relatively poor quality.
-4.2 Evaluation Metrics
-For content generation tasks, human objective evaluation is the most important evaluation method because of lacking clear ground and explicit criteria.Therefore, the human objective evaluation is the main evaluation method in our method.
-Each motion slice is re-targeted to Mixiamo model and rendered as a motion video sequence.
-They are asked to rate the slices from the following four aspects respectively: (1) Naturalness, (2) Rhythm, (3) Diversity, (4) Semantic.
+The datasets collected from Internet by video motion capture are only used for pre-training the diffusion model because of its relatively low quality.
+4.2. Evaluation Metrics
+For content generation tasks, human objective evaluation is the most important evaluation method because of lacking clear ground and explicit criteria.
+Therefore, the human objective evaluation is the main evaluation method in our experiment.
+Each motion slice is re-targeted to Mixamo [^3] https://www.mixamo.com model and rendered as a video.
+The evaluators are asked to rate the slices from the following four aspects respectively: (1) Naturalness (Nat.), (2) Rhythm (Rhy.), (3) Diversity (Div.), (4) Semantic (Sem.).
 The scores assigned to each rating are in the range of 1-10, corresponding from worst to best.
@@
 Besides, we also calculate some quantitative metrics adopted by previous works.
-FGD [49] measures the distribution difference between generated data and ground truth.
-Beat Consistency (BC) [30] calculates the average distance between every audio beat and its nearest motion beat.
+FGD (Yoon et al., 2020) measures the distribution difference between generated data and ground truth.
+Beat Consistency (BC) (Li et al., 2021b) calculates the average distance between every audio beat and its nearest motion beat.
 Diversity is a metric of measuring the variations of the synthesized gesture.
-Semantic-Relevant Gesture Recall (SRGR) [32] measures the semantic correct key point of synthesized data by comparing it with the ground truth data.
-4.3 Implementation Details
+Semantic-Relevant Gesture Recall (SRGR) (Liu et al., 2022b) measures the semantic correct key point of synthesized data by comparing it with the ground truth.
+4.3. Implementation Details
 The whole training process consists of two steps.
-In the first step, we train the VQVAE model using the Adam optimizer for 400 epochs.
-In the second step, we train the diffusion-based gesture generation model using the Adam optimizer for 3000 epochs with batchsize 256.
-The diffusion steps for training and inference are 1000.
-During the training of diffusion model, we firstly use the collected data for pre-training (2000 epochs with batchsize 256), and then use the high-quality dataset for fine-tuning (1000 epochs).
+Firstly, we train the VQVAE model by using the Adam optimizer with learning rate 0.001 for 400 epochs.
+Then, we train the diffusion-based gesture generation model using the Adam optimizer with learning rate 0.0002 for 3000 epochs with batch-size 256.
+The number of diffusion steps for training and inference is set as 1000.
+During the training of diffusion model, we firstly use the collected data for pre-training (2000 epochs with batch-size 256), and then use the high-quality dataset for fine-tuning (1000 epochs).
+In the diffusion model, the variances $\beta_{t}$ are predefined to increase linearly from 0.0001 to 0.02, with a total of 1000 noising steps set as $T=1000$.
 The length of the input audio feature is 12 seconds.
-We train the diffusion model on 32 V100 GPUs.
-4.4 Comparing Baselines
-We compare our method on BEAT dataset with the state-of-the-art methods in the last year, including DiffuseStyleGesture[44], TalkSHOW[48], QPGesture [45], EMAGE [31].
-DiffuseStyleGesture[44] is a diffusion model-based co-speech gesture synthesis approach, which takes the motion style into generation process.
-TalkSHOW[48] uses a simple full-body speech-to-motion generation framework with auto-regressive manner, in which the face, body, and hands are modeled separately.
-QPGesture [45] is a quantization-based and phase-guided motion-matching approach for co-speech gesture synthesis.
-EMAGE [31] leverages a masked gesture reconstruction to enhance audio-conditioned gesture generation.
-4.5 Evaluation Results
-Figure 3: Visualization results between the proposed method and other state-of-the-art methods.
-Qualitative Results For this task, there is lack of absolutely objective metric to evaluate the performance of various methods.
+We train the diffusion model on 32 V100 GPUs for about one week.
+4.4. Comparing Baselines
+We compare our method on BEAT dataset with several representative state-of-the-art methods in the last year, including DiffuseStyleGesture (Yang et al., 2023b), TalkSHOW (Yi et al., 2023), QPGesture (Yang et al., 2023a), EMAGE (Liu et al., 2023).
+DiffuseStyleGesture (Yang et al., 2023b) is a diffusion model-based co-speech gesture synthesis approach, which takes the motion style into generation process.
+TalkSHOW (Yi et al., 2023) uses a simple full-body speech-to-motion generation framework with auto-regressive manner, in which the face, body, and hands are modeled separately.
+QPGesture (Yang et al., 2023a) is a quantization-based and phase-guided motion-matching approach for co-speech gesture synthesis.
+EMAGE (Liu et al., 2023) leverages a masked gesture reconstruction to enhance audio-conditioned gesture generation.
+Figure 4. The visualization results of the proposed method. The gestures in the red box are semantic gesture, which is labeled by the content in green box.
+4.5. Evaluation Results
+4.5.1. Qualitative Results
+For this task, there is lack of absolutely objective metric to evaluate the performance of various methods.
 Therefore, we strongly recommend the readers to view our demonstration video to gain an intuitive understanding of the qualitative outcomes.
-In Figure 3, it shows some synthesized results of the baselines conditioned on the same speech.
+Figure 3 shows some synthesized results of these baselines conditioned on the same speech.
 From these results, we can observe that the synthesized co-speech gestures of the proposed method are more realistic, agile, and diverse than those of the baselines on BEAT datasets.
 The baselines suffer different extents of jittering in the results, especially lacking of semantic gesture and naturalness.
-DiffuseStyleGesture [44] is also a diffusion-based method, which directly learn the origin motion representations.
+DiffuseStyleGesture (Yang et al., 2023b) is also a diffusion-based method, which directly learns the origin motion representations.
 However,its results lack smoothness and naturalness (e.g.
 foot sliding).
-QPGesture [45] exhibits noticeably slower and less varied motion, while ours demonstrates agility comparable to actual movement.
+QPGesture (Yang et al., 2023a) exhibits noticeably slower and less varied motion, while ours demonstrates agility comparable to actual movement.
 For the specific semantic expressions, apart from our approach, other methods cannot effectively synthesize proper semantic gestures.
@@
 The results indicate that our method can generate proper semantic gestures for the speech semantics.
-For example, "free will" refers to the first row with open arms; "left" and "right" indicate pointing to the corresponding direction.
-These results show the surpassing capability of SIGGesture in synthesizing semantic accurate and natural co-speech gesture.
-Figure 4: The visualization results of the proposed method. The gestures in the red box are semantic gesture, which is labeled by the content in green box.
-User study.
+For example, ”free will” refers to the first row with open arms; ”left” and ”right” indicate pointing to the corresponding direction.
+These results show the surpassing capability of SIGGesture in synthesizing semantic accurate and natural gestures.
+4.5.2. User study.
 As we all known, the evaluation of generative task is very subjective.
@@
 Specifically, we randomly select 50 synthesized samples and shuffle results of different methods.
-Following previous works[52, 44], the participants are asked to scoring the gesture based on naturalness, rhythm appropriateness, diversity, and semantic consistence for the shuffled visual results..
-It’s important to note that, for diversity, we enable users to score motion diversity under the condition of smooth and natural motion.
+Following previous works(Zhi et al., 2023; Yang et al., 2023b), the participants are asked to scoring the gesture based on naturalness, rhythm appropriateness, diversity, and semantic consistence for the shuffled visual results.
+It’s important to note that, for diversity, we enable users to score motion diversity under the condition of smoothness and naturalness.
 The results of user study are shown in Table 3.
 The results show that the generated gestures of SIGGesture are dominantly preferred on four metrics over the comparing baselines.
-Especially, SIGGesture exceeds these baslines by a large margin.
+Especially, SIGGesture exceeds these baselines by a large margin in semantic relevancy.
 It can be concluded that large-scale pre-training benefits rhythm and diversity, while semantic injection is advantageous for semantic consistency.
 Overall, SIGGesture is capable of generating more realistic, synchronized, diverse, and understandable gestures that humans prefer.
-Table 3: The statistical data of user study by comparing the proposed method to baselines. The $\downarrow$" means the lower, the better. The $\uparrow$" means the higher, the better.
-[…]
-DiffuseStyleGesture (IJCAI’23) [44] 0.78 0.82 0.65 0.57
-TalkSHOW (CVPR’23)[48] 0.56 0.63 0.57 0.37
-QPGesture (CVPR’23)[45] 0.62 0.72 0.53 0.56
-EMAGE (24) [31] 0.80 0.81 0.79 0.60
+Table 3. The statistical data of user study by comparing the proposed method to baselines. The $\downarrow$” means the lower, the better. The $\uparrow$” means the higher, the better.
+Methods Nat.$\uparrow$ Rhy.$\uparrow$ Div.$\uparrow$ Sem.$\uparrow$
+DiffuseStyleGesture 0.78 0.82 0.65 0.57
+TalkSHOW 0.56 0.63 0.57 0.37
+QPGesture 0.62 0.72 0.53 0.56
+EMAGE 0.80 0.81 0.79 0.60
 SIGGesture 0.82 0.87 0.83 0.92
-Quantitative Results.
+Figure 5.
+The visual results of the proposed method on in-the-wild speeches (English, Chinese, and Japanese).
+The gestures in the red box are semantic gesture, which is labeled by the content in green box.
+4.5.3. Quantitative Results.
 We compare our method with the baselines using four metrics on BEAT dataset.
 The comparison results are shown in Table 4.
-As discussed in recent works [43, 2], these metrics have obvious weakness because these value can not keep consistent with their visual results.
-For example, TalkSHOW[48] generates unnatural body movements based on the given audio, as shown in our supplementary video, but the results of these metrics can not show this gap.
-Besides, the SRGR can not show the difference of various methods, because SRGR only considers the ground truth semantic motion for a specific speech content.
+As discussed in recent works (Tseng et al., 2023; Ao et al., 2023), these metrics have obvious weakness because these valued can not keep consistent with the visual results.
+For example, TalkSHOW (Yi et al., 2023) generates unnatural body movements based on the given audio, as shown in our supplementary video, but the results of these metrics can not show this gap.
+Besides, the SRGR can not show the difference of various methods, because SRGR only considers the ground truth semantic motion for a specific speech transcripts.
+Therefore, the SRGR of our method is lower than TalkSHOW.
+The reason is that this metric cannot accurately represent semantic performance.
+Because SRGR calculates the similarity (distance of skeleton) between the generated motion and the GT motion (with semantic labels and duration) in 3D space.
+For the same sentence, different semantic words can be expressed.
 All the same, the proposed method can outperform previous works across most metrics.
-Although we only consider two modalities (audio and speaker identity), the proposed method can excels state-of-the-art model that employs all five modalities, such as DiffStyleGesture [44].
-Table 4: The statistical data of quantitative results by comparing SIGGesture to the baselines. The $\downarrow$" means the lower, the better. The $\uparrow$" means the higher, the better.
+Although we only consider two modalities (audio and speaker identity), the proposed method can excels state-of-the-art model that employs all five modalities, such as DiffStyleGesture.
+Table 4. The statistical data of quantitative results by comparing SIGGesture to the baselines. The $\downarrow$” means the lower, the better. The $\uparrow$” means the higher, the better.
 […]
-DiffuseStyleGesture (IJCAI’23) [44] 10.14 0.233 0.504 11.975
-TalkSHOW (CVPR’23)[48] 7.313 0.279 0.463 12.859
-QPGesture (CVPR’23)[45] 19.921 0.209 0.453 9.438
-EMAGE (24) [31] 5.430 0.272 0.679 13.075
+DiffuseStyleGesture 10.14 0.233 0.504 11.975
+TalkSHOW 7.313 0.279 0.463 12.859
+QPGesture 19.921 0.209 0.453 9.438
+EMAGE 5.430 0.272 0.679 13.075
 SIGGesture 2.021 0.263 0.707 14.020
-4.6 Ablation study
+4.6. Ablation study
 In this subsection, we evaluate the contribution of two main parts, large-scale pre-training and semantic injection, by user study.
@@
 In addition to pre-training, the proposed semantic injection can significantly improve the quality of semantic gesture.
-Table 5: The statistical results of ablation study. "w/o" means "without".
+The experimental results align with our hypothesis that pre-training with large-scale data can enhance the fundamental capabilities of synthesizing rhythmic gestures, and incorporating semantic injection with LLMs can accurately generate semantic gestures.
+Table 5. The statistical results of ablation study. ”w/o” means ”without”.
 […]
@@
 SIGGesture (full) 0.82 0.87 0.83 0.92
-4.7 Generalization
-To evaluate the generalizability, we test the proposed method on a broader set of audio samples, including out-of-domain English, Chinese and Japanese.
+4.7. Generalization
+To evaluate the generalization ability, we test the proposed method on a broader set of audio samples, including out-of-domain English, Chinese and Japanese.
 The results are shown in Figure 5.
-It is also recommended to view the supplementary for more intuitive understanding and more examples.
+It is also recommended to view the supplementary video for more intuitive understanding.
 As shown in the visualization, the proposed method can accurately capture the key semantics and present precise gestures cross different language.
 The enhanced generalization capacity of our methodology is attributable to two principal components.
-Firstly, the employment of pseudo-labeled data can significantly improve the generalizability of rhythmic gestures.
+Firstly, the employment of pseudo-labeled data can significantly improve the generalization ability of rhythmic gestures.
 Secondly, disentangling the generation of rhythmic and semantic gestures by semantic injection can strengthen the controllability of semantic gesture synthesis.
 We have established an extensive collection of semantic gestures, encompassing a comprehensive range of gestural expressions.
-We leverage LLMs as a bridge to link the semantic gesture database and linguistic contexts, which has capable of robust generalization across diverse language.
+We leverage LLMs as a bridge to link the semantic gesture database with linguistic contexts, which enables robust generalization across diverse languages.
 This method effectively addresses the challenges posed by the sparsity and long-tail distribution of semantic gestures.
-As a result, the proposed method has distinguished generalizability over different speech inputs.
-Figure 5: The visual results of the proposed method on in-the-wild speeches (English, Chinese, and Japanese).
-The gestures in the red box are semantic gesture, which is labeled by the content in green box.
-5 Conclusion
-In this paper, we present a novel method named SIGGesture for high-quality co-speech gesture synthesis.
-To achieve this, we propose a semantic injection method with LLMs to improve controllability, interpretability, and generalization ability.
-With the ability of LLMs, this method can be applied to other language without any adjustment.
-Besides, we propose a large-scale dataset from Internet videos for pre-training the diffusion-based synthesis model.
-With the usage of pre-training and fine-tuning paradigm, the generated gestures are more robust to the variation of the audio.
-The proposed datasets and related statistical experiments can benefit the research of co-speech gesture synthesis, including controllable gesture synthesis, gesture foundation model in the future.
+As a result, the proposed method has distinguished generalization over different speech inputs.
+5. Conclusion
+In this paper, we introduce a novel method called SIGGesture for high-quality co-speech gesture synthesis.
+Our approach leverages a semantic injection technique using LLMs to enhance controllability, interpretability, and generalization.
+Thanks to the capabilities of LLMs, our method can be seamlessly applied to other languages without requiring any complex adjustments.
+Additionally, we have developed a large-scale 3D gesture dataset sourced from Internet videos to pre-train our diffusion-based synthesis model.
+By employing a pre-training and fine-tuning paradigm, our generated gestures exhibit greater robustness to variations in audio input.
+The proposed datasets and accompanying statistical experiments aim to advance the field of co-speech gesture synthesis, including areas such as controllable gesture synthesis and gesture foundation models.
+Looking ahead, our research will focus on generating full-body animations with enhanced and detailed expressiveness.
+6. Acknowledgements
+We thank Wenhao Ge and Baocheng Zhang for their kindly support.
+Appendix A Motion Representation
+The diffusion model synthesizes the latent feature, and then the decoder of the Vector Quantized Variational Autoencoder (VQVAE) recovers the skeletal rotation from the latent space.
+VQVAE (Oord et al., 2017) has shown excellent performance in reconstructing complex content, including image, animation, video, etc.
+This section will present the details of the VQVAE model.
+We use high-quality skeletal rotations as motion representations and parameterize these rotations using rotation matrices relative to a reference T-pose.
+The original motion is denoted as =$\mathcal{M}=\{m_{0},m_{1},...,m_{L}\}$, ${m_{i}\in R^{J\times 9}}$, where ${J}$ is the number of skeletal bones and 9 is the dimension of rotation matrix.
+The initial FPS of different datasets is converted to 30 FPS for a unified standard.
+For a specific motion sequence $\mathcal{M}$, we use VQVAE to encode and quantize it into a finite codebook $\mathcal{Z}=\{z_{i}\}_{0}^{N}$, where ${N}$ is the size of codebook and ${z_{i}}$ is a specific lexeme.
+By using VQVAE, we can quantize the continuous motion into discrete tokens marked as $\mathcal{M}_{token}=\{\mathcal{T}_{0},...,\mathcal{T}_{L/4}\}$ with 4$\times$ down-sampling.
+The framework of VQVAE is shown in Figure 6.
+A piece of original motion $\mathcal{M}$ firstly is converted to a latent embedding $\hat{e}=\{e_{0},e_{1},...,e_{L/4}\}$.
+Then, ${\hat{e}}=\{e_{0},e_{1},...,e_{L/4}\}$ is substituted by its nearest vector ${z_{j}}$ in the codebook, which can be denoted as $\arg\min\limits_{j}\parallel e_{i}-z_{j}\parallel$.
+The encoding process can be represented by the following formula
+(13) $\begin{split}\hat{e}_{i}=VQ_{encoder}(\mathcal{M}).\end{split}$
+[…]
+(14) ^$\begin{split}\hat{\mathcal{M}}=VQ_{decoder}(\hat{e}_{i}).\end{split}$
+$VQ_{decoder}$ and $VQ_{decoder}$ are modelled by Convolutional Neural Networks. In the training process, the encoder, the decoder, and the codebook are optimized by the following loss function
+(15) $\begin{split}\mathcal{L}_{VQ}=\mathcal{L}_{recon}+\parallel\hat{e}-sg(e)\parallel+\parallel sg(\hat{e})-e\parallel,\end{split}$
+where $\mathcal{L}_{recon}$ is the reconstruction error and $sg[*]$ denotes the stop gradient operator. The reconstruction error is calculated by the following formula
+[…]
+where ${\mathcal{M}^{\prime}}$ and $\mathcal{M}^{\prime\prime}$ are the velocity and acceleration of the original motion sequence $\mathcal{M}$ respectively.
+All motions are encoded into latent space for training the diffusion model.
+In our pipeline, VQVAE serves as a fundamental model used to encode motions into latent space and to recover motions from latent space.
+Figure 6. The architecture of VQVAE. It contains an encoder, a decoder and a codebook.
+Appendix B User Study
+We place significant importance on the effectiveness and fairness of our user study evaluations.
+To this end, we have recruited 30 volunteers (15 men and 15 women), including 5 researchers with expertise in gesture or animation synthesis.
+Each participant evaluates 250 co-speech gesture videos, which are generated from 50 audio clips using 5 different methods.
+To ensure that participants understand each metric and can accurately differentiate between them, we provide detailed instructions and training.
+This training covers the standard evaluation process and explains the differences between the various metrics.
+The detailed difference of these metric are as following.
+The user study focus on the following aspects, Naturalness, Rhythm, Diversity, Semantic.
+Naturalness refers to how realistically and convincingly the generated motion movements mimic real-world behavior and physics.
+This includes evaluating whether the movements are fluid, smooth, and consistent with real movements and real-world physical laws.
+Rhythm metric typically evaluates how well the visual animation aligns with the rhythm or timing of the accompanying audio.
+This involves examining whether the gesture movements (mainly are hand movements) synchronize effectively with the beats or tempo of the audio.
+Diversity metric measures how much variety is present in the generated gesture animations produced in response to a specific audio.
+A higher score of diversity ensures that the synthesized co-speech gestures are not only synchronized with the audio but also rich and varied in a range of different movements and styles rather than repeating similar patterns or showing minimal variation.
+Semantic metric evaluates how well the animations convey the intended meaning, specific semantic or emotional content of the audio.
+A high Semantic score indicates that the generated gesture accurately interprets and represents the content of the audio.
+Participants first browse the videos to familiarize themselves with the content.
+They then formally score the videos, with 5 animations from different methods displayed simultaneously on the screen.
+The differences in scores will reflect the performance variations among the different algorithms.
+Appendix C Video-based Collected dataset
+To ensure safety, raw data is collected from official TED and YiXi sources.
+Each video on TED and YiXi varies in terms of speaker’s gender, age, speech topics, and talking styles.
+This diversity enriches the dataset with a variety of gesture types.
+Finally, the dataset contains about 50% English and 50% Chinese.
+Following previous work (Yoon et al., 2019), the selection criteria for shots of interest include: containing speech content, having a stable background, displaying the entire upper body, facing the camera, and ensuring the body occupies at least 50% of the screen.
+These video clips are then processed through a gesture motion capture pipeline, depicted in Figure 7, which comprises five key components.
+Person detection utilizes Yolo-v8 (Jocher et al., 2023), while tracking employs the Sort algorithm (Bewley et al., 2016).
+(Bewley et al., 2016).
+Person key points are detected by using ViTPose (Xu et al., 2022), which are trained on a large-scale dataset.
+SMPL parameters fitting is like SMPLify (Bogo et al., 2016), which can boost the pose accuracy of SMPLX (Pavlakos et al., 2019).
+Post-processing part mainly contains motion smoothing and quality checking.
+Figure 7. The process pipeline of video-based gesture motion capture.
+Finally, the data is represented using the NEUTRAL SMPL-X skeleton. It took about 2 months and utilized 64 V100 GPUs for collecting the video-based gesture data.
+Appendix D LLM prompt
+The prompt is divided into three components: the task definition, a series of examples for few-shot learning, and the final task.
+During inference, the LLM is employed solely to generate accurate semantic gesture mappings from a predefined set.
+Our experiments utilize GPT-4 (Achiam et al., 2023) to perform this task and also involve a comparison of various prompting strategies, such as zero-shot and few-shot learning.
+In our findings, zero-shot learning prompts generally fail to produce appropriate semantic gestures.
+Conversely, few-shot learning prompts demonstrate effectiveness in performing the annotation task and yield meaningful results.
+It should be noted that these examples for few-shot learning contain both positive and negative examples.
+Below, we provide an example of the prompt, noting that some terms have been simplified due to page constraints.
+Figure 8. An example of the LLM prompt for generating proper candidate semantic gesture.